Fermi National Laboratory


Accelerator Update - May 19-May 21

Last week the Fermilab accelerators broke several records. Fermilab set a new integrated luminosity record with a level of 9159 nb-1. The Main Injector (MI) broke a record that wasn't reported, that of a high intensity of protons, 5.45E12, being sent to the Antiproton Source (Pbar) target for a single pulse.

Accelerator Overview

Monday May 19
The day shift began with Operations monitoring store 2562, and with accesses to most accelerators for maintenance.

At 9:39 AM, operators investigated a reported problem and found a small LCW leak on a Pbar Debuncher magnet (D2SF9).

Operators had beam in Linac and to the Booster dump by 6:13 PM. A Linac expert began a study period.

Operators established beam to the MI dump at 7:37 PM.

At 8:24 PM, a TeV expert began an end-of-store study with what was left of store 2562.

The Linac expert completed his studies at 8:54 PM. Operations established beam to MiniBooNe at 9:03 PM.

At 9:10 PM, the TeV suffered a sector F4 quench and lost store 2562 due to the end-of-store study.

Operations discovered that a Pbar RF station (DRF1) was having a problem. Operators managed to get some beam to circulate so they could conduct a reverse proton study at 11:03 PM. Later a number of devices had a power loss. Operators called the Duty Mechanic to investigate a possible breaker problem.

At 11:51 PM, Operations put the TeV through a dry squeeze.

Tuesday May 20
The midnight shift began with TeV, Pbar, and Recycler studies, and with MiniBooNe taking beam.

At 6:55 AM, the Linac Klystron debuncher tripped off on a water flow indication. Operators found its water pump tripped off. It reset okay, but at 11:10 AM it tripped off several more times.

Switchyard (SY) experts began attempting to run beam to the SY dump at 10:14 AM. They eventually sent beam to TeV sector F45.

MI experts set a record intensity level of 2.15E13 multi-batching at 7 PM, which is how the NUMI experiment would be getting beam when it begins operating.

At 7 PM, the TeV quenched at sector BA while Operations was trying to abort the beam. TeV experts resumed their studies at 10:13 PM.

Wednesday May 21
The midnight shift began with the TeV and Pbar conducting studies, and with MiniBooNe taking beam.

At 12:03 AM, Pbar expert completed their studies.

Pbar began stacking at 1 AM.

At 7:32 AM, MI suffered a vacuum problem that caused many vacuum valves to close. The areas must be pumped down to open the valves. This problem halted stacking.

Today's Plan
The plan for today is to conduct accelerator studies.

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/21/2003   email Fermilab